OpenStudy (anonymous):

How would one write a Maclaurin series for (x^2)e^(x^2)) in sigma notation?

OpenStudy (anonymous):

I understand the basic format, I just don't know how to deal with the \[f ^{k}(0)\] portion.

OpenStudy (kainui):

It's surprisingly easy once you know the trick, which should come as a good relief to you. Do you know the Maclauren series for \(e^x\)?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}x ^{k}/k!\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

replace \(x\) by \(x^2\) to get \(e^{x^2}\)

OpenStudy (anonymous):

But am I supposed to substitute 0 into f(x) and just insert that into the summation?

OpenStudy (anonymous):

you wrote \[e^x=\sum_{n=0}^{\infty}x ^{k}/k!\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[e^{x^2}=\sum_{n=0}^{\infty}(x^2) ^{k}/k!\]

OpenStudy (anonymous):

which you can clean up easily